TWO-PHOTON SPECTROSCOPY OF THE NO-NE COMPLEX

YANGSOO KIM, JULIAN FLENIKEN, and \underline{HENNING MEYER}, Department of Physics and Astronomy, University of Georgia, Athens, GA 30602-2451.

In our recent work on NO-Ar complexes we have shown how the rotational structure of the two-photon spectra can be analyzed in terms of the spherical tensor components of the two-photon absorption operator characteristic to the transition in the bare NO molecule.^{a,b} Due to spectral congestion we have not yet been able to analyze the NO-Ar spectra in the region of the Rydberg states F^{2}Δ and H^{2}Σ, H’^{2}Π. In this contribution, we will present the analogous results for the NO-Ne complex. Because of the weaker interaction, the various band systems correlating with the Rydberg states C^{2}Π (v_{NO}=1,2, and 4), E^{2}Σ, F ^{2}Δ, and H^{2}Σ, H’^{2}Π are well separated from each other. Apart from the H state for which a five membered stretch progression is observed, the spectra assigned to the lower Rydberg states show only transitions to two stretch vibrational levels. Spectra associated with the vibrational levels of the C^{2}Π state show very different predissociation behavior compared to the corresponding NO-Ar spectra. The results indicate important differences in electronic state mixing between the C^{2}Π and B^{2}Π states upon complexation.

aH. Meyer, J. Chem. Phys. \underline{107}, 7732(1997). bH. Meyer, J. Chem. Phys. \underline{107}, 7721(1997).